OpenStudy (anonymous):

Simplify sqrt(3 + 2 sqrt 2)

OpenStudy (anonymous):

2.4141251

OpenStudy (anonymous):

Null points...

OpenStudy (anonymous):

simplify not calculate lol

OpenStudy (anonymous):

Mind u he got a clue that way....

OpenStudy (anonymous):

oh come on what's to simplyfy in this a ninth std. student even knows how to calculate it

OpenStudy (anonymous):

I'm the tutor and I want u to simplify it...:-)

OpenStudy (anonymous):

2.4141251 Remind u of anything?

OpenStudy (anonymous):

Calculate root 2

OpenStudy (anonymous):

okay let me do it i didn't simplify it ah ...

OpenStudy (anonymous):

Let's do algebra Simplify sqrt(a+bsqrtc)

OpenStudy (anonymous):

Stop alcatraz complaining....

OpenStudy (anonymous):

ah i think I got it

OpenStudy (anonymous):

sqrt( 1 + 2 +2sqrt2) = sqrt( (1+sqrt2)^2) = 1 + sqrt2

OpenStudy (anonymous):

how's that

OpenStudy (anonymous):

oh thanks : )

OpenStudy (mimi_x3):

oh lols , ishaan pretending to be stupid xD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Come on Ishaan, solve the general case....

OpenStudy (anonymous):

ah okay

OpenStudy (anonymous):

Ramanujan was good at these..

OpenStudy (anonymous):

\[\sqrt{3+2*\sqrt{2}} = a + b*\sqrt{2}\] Square both sides: \[3+2*\sqrt{2} = (a+b*\sqrt{2})^2\] \[3+2*\sqrt{2} = a^2 + 2b^2 + 2b*\sqrt{2}\] Compare coefficients: \[2 = 2b => b = 1\] \[a^2 + 2b^2 = 3\] \[a^2 + 2 = 3\] \[a^2 = 1\] \[a = \pm1\] Therefore: \[\sqrt{2} \pm 1\] Should be correct :)

OpenStudy (anonymous):

plus only, good work.

OpenStudy (anonymous):

yep plus only, thanks :)

OpenStudy (anonymous):

by the way there was a flaw in my working which made the previous answer coincidentally right :/ Here's the correct procedure: \[\sqrt{3+2\sqrt{2}} = (a+b \sqrt{2})\] Square both sides: \[3+2 \sqrt{2} = (a+b \sqrt{2})^2\]\[3 + 2\sqrt{2} = a^2 + 2b^2 + 2ab \sqrt 2\] Comparing Coefficients: \[2ab = 2 => ab = 1\]\[a^2 + 2b^2 = 3\] Solving simultanously: \[b = {1 \over a}\] therefore: \[a^2 + 2({1 \over a^2}) - 3 = 0\] multiply both sides by a^2: \[a^4 - 3a^2 + 2 = 0\] Let x = a^2, then: \[x^2 - 3x + 2 = 0\]\[(x-1)(x-2) =0\]x = 1 or x = 2 a^2 = 1 or a^2 = 2 \[a = \pm 1 => b = \pm 1\]or\[a = \pm \sqrt{2} => b = \pm {\sqrt{2} \over 2}\] Discarding negatives, one gets the pairs \[a = 1, b = 1\] or \[a = \sqrt{2}, b = {\sqrt{2} \over 2}\] both leading to the answer \[1 + \sqrt{2}\] when substituted into \[a + b \sqrt{2}\] I apologize for the previous mistake, this is the correct working to the solution.